# On the stability of the penalty function for the $\mathbb{Z}^{2}$-hard square shift

### v2

@@ 3. Proof of the main theorem
For evaluating the difference between the dynamical sums of $\underline{x}^{i-1}$ and $\underline{x}^{i}$ over “good words”, we use an upper bound on the distance between $\sigma^{u}\underline{x}^{i-1}$ and $\sigma^{u}\underline{x}^{i}$ for each $u\in\Lambda_{N}\setminus\left(\bigcup_{n\in I^{i}_{N}}B_{n}^{i}\right)$. This upper bound depends on the distance to bad words, and this modification does not make that distance any smaller. Let $c^{i}_{n}=\lfloor\frac{\tilde{\alpha}_{n}^{i}-\tilde{\beta}_{n-1}^{i}}{2}\rfloor(\geq 0)$. Define the sets
$G_{n}^{i,+}$ $=([\tilde{\beta}_{n-1}^{i},\tilde{\beta}_{n}^{i}]\times[i,N])\setminus([\tilde{\alpha}_{n}^{i},\tilde{\beta}_{n}^{i}]\times\{i\})$
and
$G_{n}^{i,-}$ $=[\tilde{\beta}_{n-1}^{i},\tilde{\beta}_{n}^{i}]\times[-N,i-1].$
Then we obtain
$\Lambda_{N}=\bigcup_{n=\ell_{N}^{i}}^{r_{N}^{i}+1}(G_{n}^{i,+}\cup G_{n}^{i,-})\cup\bigcup_{n=\ell_{N}^{i}}^{r_{N}^{i}}[\tilde{\alpha}_{n}^{i},\tilde{\beta}_{n}^{i}].$
Now, let $0\leq i\leq N$ and consider $n\in[l_{N}^{-i},r_{N}^{-i}]$ satisfying $0\leq c_{n}^{-i}<N+i$. For $(u_{1},u_{2})\in G_{n}^{-i,+}$, the distance
$d(\sigma^{(u_{1},u_{2})}\underline{x}^{(2i-1)},\sigma^{(u_{1},u_{2})}\underline{x}^{(2i)})$
@@
is determined by three cases, and the computation is divided into four regions:
$A$ =$=\{(u_{1},u_{2}):\tilde{\beta}^{-i}_{n-1}+1\leq u_{1}\leq\tilde{\beta}^{-i}_{n-1}+\tilde{c}_{n}^{-i},-i\leq u_{2}\leq-i+u_{1}-(\tilde{\beta}_{n-1}^{-i}+1)\}$
$\quad\cup\{(u_{1},u_{2}):\tilde{\beta}^{-i}_{n-1}+\tilde{c}_{n}^{-i}+1\leq u_{1}\leq\tilde{\alpha}_{n}^{-i}-1,-i\leq u_{2}\leq-i+\tilde{c}_{n}^{-i}-1-u_{1}+(\tilde{\beta}_{n-1}^{-i}+\tilde{c}_{n}^{-i}+1)\};$
$B$ =$=\{(u_{1},u_{2}):\tilde{\beta}^{-i}_{n-1}+1\leq u_{1}\leq\tilde{\beta}^{-i}_{n-1}+\tilde{c}_{n}^{-i},-i+u_{1}-(\tilde{\beta}_{n-1}^{-i}+1)\leq u_{2}\leq-i+\tilde{c}_{n}^{-i}-1\}$
$\quad\cup\{(u_{1},u_{2}):\tilde{\beta}^{-i}_{n-1}+\tilde{c}_{n}^{-i}+1\leq u_{1}\leq\tilde{\alpha}_{n}^{-i}-1,-i+u_{1}-(\tilde{\alpha}_{n}^{-i}-1)\leq u_{2}\leq-i+\tilde{c}_{n}^{-i}-1\};$
$C$ =$=\{(u_{1},u_{2}):\tilde{\alpha}_{n}^{-i}\leq u_{1}\leq\tilde{\beta}_{n}^{-i},-i+1\leq u_{2}\leq\tilde{c}_{n}^{-i}\};$
$D$ $=[\tilde{\beta}_{n-1}^{-i}+1,\tilde{\beta}_{n}^{-i}-1]\times[-i+\tilde{c}_{n}^{-i}+1,N].$
@@
$d(\sigma^{(u_{1},u_{2})}\underline{x}^{(2i-1)},\sigma^{(u_{1},u_{2})}\underline{x}^{(2i)})=\left\{\begin{array}[]{ccc}\frac{1}{2^{u_{1}-\tilde{\beta}_{n-1}^{-i}}}&\mbox{if}&u_{1}\leq\tilde{\beta}_{n-1}^{-i}+c_{n}^{-i}\\ \frac{1}{2^{\tilde{\alpha}_{n}^{-i}-u_{1}}}&\mbox{if}&u_{1}>\tilde{\beta}_{n-1}^{-i}+c_{n}^{-i}.\end{array}\right.$
@@
$d(\sigma^{(u_{1},u_{2})}\underline{x}^{(2i-1)},\sigma^{(u_{1},u_{2})}\underline{x}^{(2i)})=\frac{1}{2^{u_{2}}}.$
@@
$S_{G_{n}^{-i,+}}g(\underline{x}^{(2i-1)})-S_{G_{n}^{-i,+}}g(\underline{x}^{(2i)})$ $<\left(\sum_{k=1}^{c_{n}^{i}}2k\cdot\frac{1}{2^{k}}+\sum_{k=1}^{c_{n}^{i}}2k\cdot\frac{1}{2^{k}}+\sum_{k=1}^{c_{n}^{i}}(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\frac{1}{2^{k}}\right.$
$\hskip 30.00005pt\left.+\sum_{k=c_{n}^{i}+1}^{N-i}(\tilde{\beta}^{i}_{n}-\tilde{\beta}^{i}_{n-1})\frac{1}{2^{k}}+\sum_{k=1}^{c_{n}^{i}}\frac{2}{2^{k}}\right)\varepsilon$
$<\left(\sum_{k=1}^{c_{n}^{i}}\frac{4k+2}{2^{k}}+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\sum_{k=1}^{N-i}\frac{1}{2^{k}}+\sum_{k=c_{n}^{i}+1}^{N-i}\frac{(\tilde{\alpha}^{i}_{n}-\tilde{\beta}_{n-1})}{2^{k}}\right)\varepsilon$
$\hskip 44.0pt(\because\tilde{\beta}_{n}^{i}-\tilde{\beta}_{n-1}^{i}=\tilde{\beta}_{n}^{i}-\tilde{\alpha}_{n}^{i}+\tilde{\alpha}_{n}^{i}-\tilde{\beta}_{n-1}^{i})$
$\leq\left(\sum_{k=1}^{c_{n}^{i}}\frac{4k+2}{2^{k}}+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\sum_{k=1}^{N-i}\frac{1}{2^{k}}+\sum_{k=c_{n}^{i}+1}^{N-i}\frac{2k}{2^{k}}\right)\varepsilon$
$\hskip 44.0pt(\because\tilde{\alpha}^{i}_{n}-\tilde{\beta}_{n-1}\leq 2c_{n}^{i}+1)$
$\leq\left(\sum_{k=1}^{N-i}\frac{4k+2}{2^{k}}+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\sum_{k=1}^{N-i}\frac{1}{2^{k}}\right)\varepsilon$
(4) $=\left(14+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\right)\varepsilon.$
@@
For $n\in[\ell_{N}^{-i},r_{N}^{-i}]$ such that $N\leq c_{n}^{-i}\leq N+i$ there are no regions D and A in Figure2, and region B is cut off in the middle. Hence it is easy to see that we have
$S_{G_{n}^{-i,-}}g(\underline{x}^{(2i-1)})-S_{G_{n}^{-i,-}}g(\underline{x}^{(2i)})$ $<(14+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n}))\varepsilon.$
By the same argument for $10\leq i\leq N-$ 1 and $n\in[r_{N}^{i},\ell_{N}^{i}]$ we have
(5) $S_{G_{n}^{i,\pm}}g(\underline{x}^{2i})-S_{G_{n}^{i,\pm}}g(\underline{x}^{2i+1})$ $<(14+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})).$
Noting that $G_{n}^{-i,+}\cap G_{n}^{-i,+}=\emptyset$, we have
$\sum_{i=0}^{2N}\left(S_{\Lambda_{N}}g(\underline{x}^{i-1})-S_{\Lambda_{N}}g(\underline{x}^{i})\right)$ $\leq\sum_{i=0}^{N}\sum_{n=\ell_{N}^{-i}}^{r_{N}^{-i}}\left(S_{G_{n}^{-i,+}\cup G_{n}^{-i,-}}g(\underline{x}^{(2i-1)})-S_{G_{n}^{-i,+}\cup G_{n}^{-i,-}}g(\underline{x}^{(2i)})\right)$
$\quad+\sum_{i=0}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\left(S_{G_{n}^{i,+}\cup G_{n}^{i,-}}g(\underline{x}^{(2i)})-S_{G_{n}^{i,+}\cup G_{n}^{i,-}}g(\underline{x}^{(2i+1)})\right)$
$\quad\quad+\sum_{i=0}^{N}\sum_{n=\ell_{N}^{-i}}^{r_{N}^{-i}}\left(S_{B_{n}^{-i}\times\{-i\}}g(\underline{x}^{(2i-1)})-S_{B_{n}^{-i}\times\{-i\}}g(\underline{x}^{(2i)})\right)$
$\quad\quad\quad+\sum_{i=0}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\left(S_{B_{n}^{i}\times\{i\}}g(\underline{x}^{(2i)})-S_{B_{n}^{i}\times\{i\}}g(\underline{x}^{(2i+1)})\right)$
$<2\sum_{i=-N}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}(14+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n}))\varepsilon+\sum_{i=-N}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\#B_{n}^{i}(-1+2\varepsilon)$
$\leq 28\varepsilon\sum_{i=-N}^{N}\#I_{N}^{i}+2\varepsilon\sum_{i=-N}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\#B_{n}^{i}+\sum_{i=-N}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\#B_{n}^{i}(-1+2\varepsilon)$
$\leq(28\varepsilon+2\varepsilon-1+2\varepsilon)\#\{u\in\Lambda_{N}:\sigma^{u}\underline{x}\in I^{c}\}$
$=(-1+32\varepsilon)\#\{u\in\Lambda_{N}:\sigma^{u}\underline{x}\in I^{c}\}.$
Dividing by $(2N+1)^{2}$ by the both sides of (2), we have
$\frac{1}{(2N+1)^{2}}\left(S_{\Lambda_{N}}g(\underline{x})-S_{\Lambda_{N}}g(\tilde{\underline{x}})\right)$ $<\frac{2}{(2N+1)}+\frac{1}{(2N+1)^{2}}(-1+32\varepsilon)\#\{u\in\Lambda_{N}:\sigma^{u}\underline{x}\in I^{c}\}$
$\leq\frac{2}{(2N+1)}-\frac{1}{2}\frac{1}{(2N+1)^{2}}\#\{u\in\Lambda_{N}:\sigma^{u}\underline{x}\in I^{c}\}.$
@@
, which complete the proof.

### v3

@@ 3. Proof of the main theorem
For evaluating the difference between the Birkhoff sums of $\underline{x}^{(i-1)}$ and $\underline{x}^{(i)}$ over “unchanged words,” we use an upper bound on the distance between $\sigma^{\bm{u}}\underline{x}^{(i-1)}$ and $\sigma^{\bm{u}}\underline{x}^{(i)}$ for each $\bm{u}\in\Lambda_{N}\setminus S^{i}$. By Lemma 3.1, this upper bound depends on the distance to bad words when $f(\sigma^{\bm{u}}\underline{x}^{(i-1)})=f(\sigma^{\bm{u}}\underline{x}^{(i)})$.
By the definition of the penalty function, we have $f(\sigma^{\bm{u}}\underline{x}^{(i-1)})\neq f(\sigma^{\bm{u}}\underline{x}^{(i)})$ if $\bm{u}\in S^{i}$ or $\bm{u}=\bm{v}-e_{i}$ for some $\bm{v}\in S^{i}$ and $i=1,2$.
In the latter case, we have
$g$ $(\sigma^{\bm{u}}\underline{x}^{(i-1)})-g(\sigma^{\bm{u}}\underline{x}^{(i)})$
$\quad=g(\sigma^{\bm{u}}\underline{x}^{(i-1)})-f(\sigma^{\bm{u}}\underline{x}^{(i-1)})+f(\sigma^{\bm{u}}\underline{x}^{(i-1)})-f(\sigma^{\bm{u}}\underline{x}^{(i)})+f(\sigma^{\bm{u}}\underline{x}^{(i)})-g(\sigma^{\bm{u}}\underline{x}^{(i)}).$
Since the change from $\underline{x}^{(i-1)}$ to $\underline{x}^{(i)}$ does not introduce any new bad words, we obtain
$g(\sigma^{\bm{u}}\underline{x}^{(i-1)})-g(\sigma^{\bm{u}}\underline{x}^{(i)})\leq\varepsilon\,d(\sigma^{\bm{u}}\underline{x}^{(i-1)},\sigma^{\bm{u}}\underline{x}^{(i)}).$
First we consider $R^{t(i)}$ and let $\beta_{0}^{t(i)}=-N-1$ and $\alpha_{K_{t(i)}+1}^{t(i)}=\beta_{K_{t(i)}+1}^{t(i)}=N+1$. Then for each $1\leq k\leq K_{t(i)}$, set $c^{t(i)}_{k}=\lfloor\frac{\alpha_{k}^{t(i)}-\beta_{k-1}^{t(i)}}{2}\rfloor(>0)$ and the sets
$G_{k}^{t(i)}$ $=([\beta_{k-1}^{t(i)},\beta_{k}^{t(i)}]\times[i,N])\setminus([\alpha_{k}^{t(i)},\beta_{k}^{t(i)}]\times\{i\})$
$=([\beta_{k-1}^{t(i)},\beta_{k}^{t(i)}]\times[i,N])\setminus B_{k}^{t(i)}.$
Remark that we have
$\bigcup_{k=1}^{K_{t(i)}+1}G_{k}^{t(i)}$ $=[-N,N]\times[i,N]\setminus\left(\bigcup_{k=1}^{K_{t(i)}}B_{k}^{t(i)}\right)$
$=R^{t(i)}\setminus\left(\left(\bigcup_{k=1}^{K_{t(i)}}B_{k}^{t(i)}\right)\cup\{(u_{1},u_{2})\mid u_{1}\in[-i,i],|u_{1}|\leq u_{2}\leq i-1\}\right).$
@@
Take $1\leq k\leq K_{t(k)}$ such that $c^{t(i)}_{k}<N-i$. For $\bm{u}\in G_{k}^{t(i)}$, the distance $d(\sigma^{\bm{u}}\underline{x}^{(i-1)},\sigma^{\bm{u}}\underline{x}^{(i)})$ is determined by three cases, and the computation is divided into four regions:
$A$ =$=\{(u_{1},u_{2})\mid\beta^{t(i)}_{k-1}+1\leq u_{1}\leq\beta^{t(i)}_{k-1}+c_{k}^{t(i)},-i\leq u_{2}\leq-i+u_{1}-(\beta^{t(i)}_{k-1}+1)\}$
$\quad\cup\{(u_{1},u_{2})\mid\beta^{t(i)}_{k-1}+c_{k}^{t(i)}+1\leq u_{1}\leq\alpha_{k}^{t(i)}-1,-i\leq u_{2}\leq-i+c_{k}^{t(i)}-1-u_{1}+(\beta^{t(i)}_{k-1}+c_{k}^{t(i)}+1)\};$
$B$ =$=\{(u_{1},u_{2})\mid\beta^{t(i)}_{k-1}+1\leq u_{1}\leq\beta^{t(i)}_{k-1}+c_{k}^{t(i)},-i+u_{1}-(\beta^{t(i)}_{k-1}+1)\leq u_{2}\leq-i+c_{k}^{t(i)}-1\}$
$\quad\cup\{(u_{1},u_{2})\mid\beta^{t(i)}_{k-1}+c_{k}^{t(i)}+1\leq u_{1}\leq\alpha_{k}^{t(i)}-1,-i+u_{1}-(\alpha_{k}^{t(i)}-1)\leq u_{2}\leq-i+c_{k}^{t(i)}-1\};$
$C$ =$=\{(u_{1},u_{2})\mid\alpha_{k}^{t(i)}\leq u_{1}\leq\beta_{k}^{t(i)},-i+1\leq u_{2}\leq c_{k}^{t(i)}\};$
$D$ $=[\beta^{t(i)}_{k-1}+1,\beta_{k}^{t(i)}-1]\times[-i+c_{k}^{t(i)}+1,N].$
@@
$d(\sigma^{(u_{1},u_{2})}\underline{x}^{(i-1)},\sigma^{(u_{1},u_{2})}\underline{x}^{(i)})=\left\{\begin{array}[]{ccc}\frac{1}{2^{u_{1}-\beta_{k-1}^{t(i)}}}&\mbox{if}&u_{1}\leq\beta_{k-1}^{t(i)}+c_{k}^{t(i)}\\ \frac{1}{2^{\tilde{\alpha}_{n}^{-i}-u_{1}}}&\mbox{if}&u_{1}>\beta_{k-1}^{t(i)}+c_{k}^{t(i)}.\end{array}\right.$
@@
$d(\sigma^{(u_{1},u_{2})}\underline{x}^{(i-1)},\sigma^{(u_{1},u_{2})}\underline{x}^{(i)})=\frac{1}{2^{u_{2}}}.$
@@
$S_{G_{k}^{t(i)}}g(\underline{x}^{(i-1)})-S_{G_{k}^{t(i)}}g(\underline{x}^{(i)})$ $<\left(\sum_{\ell=1}^{c_{k}^{t(i)}}2\ell\cdot\frac{1}{2^{\ell}}+\sum_{\ell=1}^{c_{k}^{t(i)}}2\ell\cdot\frac{1}{2^{\ell}}+\sum_{\ell=1}^{c_{k}^{t(i)}}(\beta^{t(i)}_{k}-\alpha^{t(i)}_{k})\frac{1}{2^{\ell}}\right.$
$\hskip 30.00005pt\left.+\sum_{\ell=c_{k}^{t(i)}}^{N-i}(\beta^{t(i)}_{k}-\beta^{t(i)}_{k-1})\frac{1}{2^{\ell}}+\sum_{\ell=1}^{c_{k}^{t(i)}}\frac{2}{2^{\ell}}\right)\varepsilon$
$<\left(\sum_{\ell=1}^{c_{k}^{t(i)}}\frac{4\ell+2}{2^{\ell}}+(\beta^{t(i)}_{k}-\alpha_{k}^{t(i)})\sum_{\ell=1}^{N-i}\frac{1}{2^{\ell}}+\sum_{\ell=c_{k}^{t(i)}+1}^{N-i}\frac{(\alpha_{k}^{t(i)}-\beta^{t(i)}_{k-1})}{2^{\ell}}\right)\varepsilon$
$\hskip 44.0pt(\because\beta^{t(i)}_{k}-\beta^{t(i)}_{k-1}=\beta^{t(i)}_{k}-\alpha_{k}^{t(i)}+\alpha_{k}^{t(i)}-\beta^{t(i)}_{k-1})$
$\leq\left(\sum_{\ell=1}^{c_{k}^{t(i)}}\frac{4\ell+2}{2^{\ell}}+(\beta^{t(i)}_{k}-\alpha_{k}^{t(i)})\sum_{\ell=1}^{N-i}\frac{1}{2^{\ell}}+\sum_{\ell=c_{k}^{t(i)}+1}^{N-i}\frac{2\ell}{2^{\ell}}\right)\varepsilon$
$\hskip 44.0pt(\because\alpha_{k}^{t(i)}-\beta^{t(i)}_{k-1}\leq 2c_{k}^{t(i)}+1)$
$\leq\left(\sum_{\ell=1}^{N-i}\frac{4\ell+2}{2^{\ell}}+(\beta^{t(i)}_{k}-\alpha_{k}^{t(i)})\sum_{\ell=1}^{N-i}\frac{1}{2^{\ell}}\right)\varepsilon$
(3) $=\left(14+(\beta^{t(i)}_{k}-\alpha_{k}^{t(i)})\right)\varepsilon.$
@@
For $1\leq k\leq K^{t(i)}$ such that $N-i\leq c_{k}^{t(i)}$ there are no regions D and A in Figure 3, and region B is cut off in the middle. Hence it is easy to see that we have
$S_{G_{k}^{t(i)}}g(\underline{x}^{(i-1)})-S_{G_{k}^{t(i)}}g(\underline{x}^{(i)})$ $<(14+(\beta^{t(i)}_{k}-\alpha^{t(i)}_{k}))\varepsilon.$
Then we have
$S_{[-N,N]\times[i,N]}g(\underline{x}^{(i)})-S_{[-N,N]\times[i,N]}g(\underline{x}^{(i-1)})$ $<\sum_{k=1}^{K_{t(i)}+1}(14+(\beta_{k}^{t(i)}-\alpha_{k}^{t(i)}))\varepsilon+\sum_{k=1}^{K_{t(i)}}|B_{k}^{t(i)}|(-1+2\varepsilon)$
$\leq 14(K_{t(i)}+1)\varepsilon+\sum_{k=1}^{K_{t(i)}}|B_{k}^{t(i)}|(-1+3\varepsilon)$
$\leq 14\varepsilon+(-1+17\varepsilon)|S^{t(i)}|,$
where the last inequality holds because $K_{t(i)}\leq|S^{t(i)}|$.
By the similar argument, the estimate on the remain region $\{(u_{1},u_{2})\mid u_{1}\in[-i,i],|u_{1}|\leq u_{2}\leq i-1\}=R^{t(i)}\cap\Lambda_{i-1}$ is bounded by
$S_{R^{t(i)}\cap\Lambda_{i-1}}g(\underline{x}^{i})-S_{R^{t(i)}\cap\Lambda_{i-1}}g(\underline{x}^{i-1})$ $<\sum_{k=1}^{K_{t(i)}+1}(14+(\beta_{k}^{t(i)}-\alpha_{k}^{t(i)}))\varepsilon$
$\leq 14\varepsilon+15|S^{t(i)}|\varepsilon$
Then we have
$S_{R^{t(i)}}g(\underline{x}^{(i)})-S_{R^{t(i)}}g(\underline{x}^{i-1})<28\varepsilon+(-1+32\varepsilon)|S^{t(i)}|.$
For $R^{(b(t))},R^{(r(i))},R^{l(i)}$, by the similar argument we have
$S_{R^{\tau}}g(\underline{x}^{(i)})-S_{R^{\tau}}g(\underline{x}^{i-1})<28\varepsilon+(-1+32\varepsilon)|S^{t(i)}|.$
for $\tau\in\{b(i),r(i),l(i)\}$. Combining all, we have
$S_{\Lambda_{N}}g(\underline{x}^{(i)})-S_{\Lambda_{N}}g(\underline{x^{(i-1)}})<112\varepsilon+(-1+32\varepsilon)\sum_{\tau}|S^{\tau}|=112\varepsilon+(-1+32\varepsilon)|S^{i}|.$
Dividing the both sides of (1) by $(2N+1)^{2}$,, we have
$\frac{1}{(2N+1)^{2}}(S_{\Lambda_{N}}g(\underline{x})-S_{\Lambda_{N}}g(\tilde{\underline{x}}))$ $\leq\frac{1}{(2N+1)^{2}}\sum_{i=0}^{N}(112\varepsilon+(-1+32\varepsilon)|S^{i}|)+\frac{2(8N+1)}{(2N+1)^{2}}$
$\leq\frac{112\varepsilon(NL1)+2(8N+1)}{(2N+1)^{2}}-\frac{1}{2}\frac{1}{(2N+1)^{2}}\sum_{i=0}^{N}|S^{i}|$
$=\frac{112\varepsilon(N+1)+2(8N+1)}{(2N+1)^{2}}-\frac{1}{2}\frac{1}{(2N+1)^{2}}\#\{\bm{u}\in\Lambda_{N}\mid\sigma^{\bm{u}}\underline{x}\in I^{c}\}$
@@
which complete the proof. ∎